### v4

# Unexpected properties of Markov chains with an infinitely countable set of states

Abstract.
For a class of irreducible Markov chains with an infinitely countable set of states, we establish a new verifiable necessary and sufficient condition for recurrence and transience. We show that if one of the basic assumptions is not satisfied, then the statement of the theorem becomes invalid.
Key words and phrases:
Discrete-time Markov chains; recurrence and transience; birth-and-death processes; continuous-time Markov processes, stochastic calculus
1991 Mathematics Subject Classification:
60J10, 60J80, 60J27, 60H07
Corresponding author
1. Introduction
Let $\mathscr{P}$ be an irreducible Markov chain with the infinite set of states $\{0,1,\ldots\}$. Let $p_{i,j}^{(n)}$ denote $n$-step transition probabilities from state $i$ to state $j$, $p_{i,j}=p_{i,j}^{(1)}$. Recall that a Markov chain is recurrent, if its initial state 00 is visited infinitely many times with probability 1. It is well-known that the Markov chain $\mathscr{P}$ is recurrent, if and only if $\sum_{n=1}^{\infty}p_{0,0}^{(n)}=\infty$ (see [8]). The application of this last condition is very limited, since the derivation of $p_{0,0}^{(n)}$ is generally combinatorially hard. The known applications are chiefly related to simple random walks and can be found in many textbooks on the theory of Markov processes.
In this paper, we suggest a new result that enables us to verify the condition of recurrence or transience for a relatively wide class of irreducible Markov chains. It turns out that the result presented here is an extension of the well-known one from the theory of birth-and-death processes, but to our best knowledge it is unknown. The proof of the main result is based on the reduction to a continuous time Markov process, and, as a result, we arrive at the equations that describe a birth-and-death process. Then the well-known results of the birth-and-death theory are applied. The suggested approach makes the proof of the main result simpler compared to possible routine analysis that would use direct analytic derivations.
Definition 1.1.
The states of Markov chain $\mathscr{P}$ are said to form connected domain if for any $i\geq 1$ there are $j_{1}(i)$ and $j_{2}(i)$, $0\leq j_{1}(i)<i<j_{2}(i)\leq\infty$ such that $p_{i,k}>0$ for all $j_{1}(i)\leq k\leq i-1$ and $i+1\leq k\leq j_{2}(i)$, and $\sum_{k=j_{1}(i)}^{j_{2}(i)}p_{i,k}=1$.
An example of Markov chain, the states of which form connected domain is a Markov chain, where $p_{i,j}>0$ for all $i,j\geq 1$ with $i\neq j$. The examples of Markov chains the states of which do not form connected domain are considered in Section 3.
Theorem 1.2.
Assume that the states of the Markov chain $\mathscr{P}$ form connected domain, and $\sum_{j=1}^{\infty}jp_{0,j}<\infty$. Denote $e_{i}^{-}=\sum_{j=1}^{\infty}jp_{i+j-1,i-1}$ and $e_{i}^{+}=\sum_{j=1}^{i+1}jp_{i-j+1,i+1}$, $i\geq 1$, assuming that $e_{i}^{-}<\infty$.Then, the Markov chain $\mathscr{P}$ is recurrent if and only if
$\sum_{n=1}^{\infty}\prod_{i=1}^{n}\frac{e_{i}^{-}}{e_{i}^{+}}=\infty.$
Remark 1.3.
The assumptions $\sum_{j=1}^{\infty}jp_{0,j}<\infty$ and $e_{i}^{-}<\infty$, $i\geq 1$, are technical, since the infinite series $\sum_{j=1}^{\infty}jp_{0,j}$ and $\sum_{j=1}^{\infty}jp_{i+j-1,i-1}$ appears explicitly in the proof of the theorem. It is possible to avoid these assumptions using additional arguments. Specifically, one can consider the series of Markov chains, in which the required parameters are assumed to increase to infinity and arrive at the conclusion in the limit case.
The rest of this paper is organized as follows. In Section 2, we prove the theorem. In Section 3, we provide a nontrivial example of recurrent Markov chain demonstrating how the statement of the theorem works. In Section 4, we provide counterexamples to demonstrate that the key assumption of the theorem, that the states of a Markov chain forms connected domain, is important. In Section 5, we discuss our findings.
2. Proof of the theorem
We start the proof from the preliminary information given in Sections 2.1 and 2.2. In section 2.1, we recall information about Markov chains, the transition probability matrix of which describes a birth-and-death process. In Section 2.2, some natural extension of those Markov chains helping us to further simplify the proof is provided. In Section 2.3 the main part of the proof is given. We provide a number of illustrative examples that clarify the further steps of the proof.
2.1. Birth-and-death process
The important class of Markov chains we deal with in this paper is the Markov chain $\mathscr{X}$ that is described by the following transition probability matrix:
(1) $\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ q_{1}&0&p_{1}&0&0&0&\ldots\\ 0&q_{2}&0&p_{2}&0&0&\ldots\\ 0&0&q_{3}&0&p_{3}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right),\quad q_{i}+p_{i}=1,\quad i\geq 1.$
It describes transition probabilities for a birth-and-death process. It is well-known that the Markov chain $\mathscr{X}$ is recurrent if and only if $\sum_{n=1}^{\infty}\prod_{i=1}^{n}q_{i}/p_{i}=\infty$ (see [6], [4, p. 286]).
2.2. Simplification of the problem
We start from the following notice related to natural extension of (1). Let $c_{1}$, $c_{2}$,…be a sequence, $0<c_{i}\leq 1$, $i\geq 1$. Consider the Markov chain $\mathscr{Y}$ with the transition probabilities matrix
(2) $\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ c_{1}q_{1}&1-c_{1}&c_{1}p_{1}&0&0&0&\ldots\\ 0&c_{2}q_{2}&1-c_{2}&c_{2}p_{2}&0&0&\ldots\\ 0&0&c_{3}q_{3}&1-c_{3}&c_{3}p_{3}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right).$
We prove first that the Markov chain $\mathscr{Y}$ is recurrent if and only if the Markov chain $\mathscr{X}$ is recurrent. Let $x_{i,j}$ denote the transition probability from state $i$ to state $j$ of the Markov chain $\mathscr{X}$, and let $y_{i,j}$ denote the transition probability from state $i$ to state $j$ of the Markov chain $\mathscr{Y}$. Apparently, that $x_{i,i-1}$ and $x_{i,i+1}$ can be considered as the conditional transition probabilities in the Markov chain $\mathscr{Y}$ given that the state $i$ of this Markov chain changes. Therefore, if the initial state of the Markov chain $\mathscr{X}$ is visited only finite number of times with probability 1, the same is true for the Markov chain $\mathscr{Y}$. Furthermore, the number of visits state 00 for both of the Markov chains $\mathscr{X}$ and $\mathscr{Y}$ has the same distribution. Otherwise, if the initial state of the Markov chain $\mathscr{X}$ is visited infinitely many number of times with probability 1, then the same is true for the Markov chain $\mathscr{Y}$ and vice versa.
Based on this specific example, one can assume without loss of generality that $p_{i,i}=0$ for all $i$, and under this additional assumption we shall prove our main result.
2.3. Chapman-Kolmogorov equations
We show that the structure of the equations for the final probabilities can be reduced to the same structure as that of birth-and-death processes. For this purpose, it is convenient to switch from the original Markov chain to the continuous time Markov process as follows. We assume that the times between transitions all are independent and exponentially distributed with mean 1.
First, we give three examples to make our further arguments clearer. The first two examples describe elementary models. The third one describes a natural model and is closer to the formulated result. The further example of Section 3 is based on the model considered in the third example of this section.
Example 2.1.
Let $\mathscr{A}$ be a birth-and-death process, where the birth and death rates are $\lambda_{n}>0$ and $\mu_{n}>0$ when there are $n$ particles in the population. Let $\mathscr{B}$ be another stochastic process with the following behavior. When there are $n$ particles in the population, there can occur either two births simultaneously (i.e. the birth of twins) or one death. The birth rate is assumed to be $\lambda_{n}/2$ (i.e. times between births of twins are exponentially distributed with mean $2/\lambda_{n}$) while the death rate is $\mu_{n}$. Let $\lambda_{0}$ be an arbitrary positive regeneration parameter for the birth-and-death process $\mathscr{A}$ and let $\lambda_{0}/2$ be such one for the process $\mathscr{B}$, when there are no individuals in these populations. According to the traditional definition, the stochastic process $\mathscr{B}$ is not a birth-and-death process. However, for our convenience it will be further called birth-and-death process as well. It will be shown in this example that the birth-and-death process $\mathscr{B}$ is recurrent if and only if the birth-and-death process $\mathscr{A}$ is recurrent.
The well-known system of equations for birth-and-death processes that describes the process $\mathscr{A}$ is
(3) $P_{n}\left(\lambda_{n}+\mu_{n}\right)$ $=$ $\mu_{n+1}P_{n+1}+\lambda_{n-1}P_{n-1},\ n\geq 1,$
(4) $\lambda_{0}P_{0}$ $=$ $\mu_{1}P_{1},$
where $P_{0},P_{1},\ldots$ denote the final probabilities (if exist).
For the process $\mathscr{B}$, we will derive the system of equation
(5) $Q_{n}(\lambda_{n}+\mu_{n})$ $=$ $\mu_{n+1}Q_{n+1}+\lambda_{n-2}Q_{n-2},\ n\geq 2,$
(6) $Q_{1}(\lambda_{1}+\mu_{1})$ $=$ $\mu_{2}Q_{2},$
(7) $\lambda_{0}Q_{0}$ $=$ $\mu_{1}Q_{1},$
where $Q_{0},Q_{1},\ldots$ denote the final probabilities (if exist). The system of equations (5)–(7) cannot be derived by the standard application of Chapman-Kolmogorov equations, since the fact that the up-crossing jump is not equal to $1$ adds certain complexities for justification of these equations. Therefore, our derivation is based on stochastic calculus, that helps us to resolve the issue and easily justify the system of equations (5)–(7) (see also [3] for the solution of a closely related problem). Our approach is then implicitly used for the more complex systems considered in the sequel. (The proofs are not repeated because of their similarity.)
For the birth-and-death process $\mathscr{B}$, let càdlàg process $Z(t)$ denote the number of individuals in the population in time $t$, $Z(0)=0$.
For $n\geq 2$ we have
(8) $\mathsf{I}\{Z(t)=n\}=\mathsf{I}\{Z(t-)=n+1\}\mathsf{I}\{\mathrm{M}_{n+1}(t)-\mathrm{M}_{n+1}(t-)=1\}$
$+\mathsf{I}\{Z(t-)=n-2\}\mathsf{I}\{\tilde{\Lambda}_{n-2}(t)-\tilde{\Lambda}_{n-2}(t-)=2\}$
$+\mathsf{I}\{Z(t-)=n\}\mathsf{I}\{\mathrm{M}_{n}(t)-\mathrm{M}_{n}(t-)=0\}\mathsf{I}\{\tilde{\Lambda}_{n}(t)-\tilde{\Lambda}_{n}(t-)=0\},$
where $\mathrm{M}_{n}(t)$ denotes the Poisson process with rate $\mu_{n}$, $\tilde{\Lambda}_{n}(t)=2\Lambda_{n}(t)$, $\Lambda_{n}(t)$ denotes the Poisson process with rate $\lambda_{n}/2$.
For $n=0,1$, we have
(9) $\mathsf{I}\{Z(t)=1\}=\mathsf{I}\{Z(t-)=2\}\mathsf{I}\{\mathrm{M}_{2}(t)-\mathrm{M}_{2}(t-)=1\}$
$+\mathsf{I}\{Z(t-)=1\}\mathsf{I}\{\mathrm{M}_{1}(t)-\mathrm{M}_{1}(t-)=0\}\mathsf{I}\{\tilde{\Lambda}_{1}(t)-\tilde{\Lambda}_{1}(t-)=0\},$
(10) $\mathsf{I}\{Z(t)=0\}=\mathsf{I}\{Z(t-)=1\}\mathsf{I}\{\mathrm{M}_{1}(t)-\mathrm{M}_{1}(t-)=1\}$
$+\mathsf{I}\{Z(t-)=0\}\mathsf{I}\{\tilde{\Lambda}_{0}(t)-\tilde{\Lambda}_{0}(t-)=0\}.$
From (8), (9) and (10) we arrive at the system of stochastic differential equations:
(11) $\mathrm{d}\mathsf{I}\{Z(t)=n\}=$ $\mathsf{I}\{Z(t-)=n+1\}\mathrm{d}\mathrm{M}_{n+1}(t)$
$+2\mathsf{I}\{Z(t-)=n-2\}\mathrm{d}\Lambda_{n-2}(t)$
$-\mathsf{I}\{Z(t-)=n\}(\mathrm{d}\mathrm{M}_{n}(t)+2\mathrm{d}\Lambda_{n}(t)),\ n\geq 2,$
$\mathrm{d}\mathsf{I}\{Z(t)=1\}=$ $\mathsf{I}\{Z(t-)=2\}\mathrm{d}\mathrm{M}_{2}(t)$
$-\mathsf{I}\{Z(t-)=1\}(\mathrm{d}\mathrm{M}_{1}(t)+2\mathrm{d}\Lambda_{1}(t)),$
$\mathrm{d}\mathsf{I}\{Z(t)=0\}=\mathsf{I}\{Z(t-)=1\}\mathrm{d}\mathrm{M}_{1}(t)-\mathsf{I}\{Z(t-)=0\}\mathrm{d}\Lambda_{0}(t).$
Let us now derive (5) from (11). The derivation of (6) and (7) is similar and therefore not considered here. For the Poisson processes $\mathrm{M}_{n}(t)$ and $\Lambda_{n}(t)$ we use semimartingale decompositions $\mathrm{M}_{n}(t)=\mu_{n}t+m_{\mathrm{M}_{n}}(t)$ and $\Lambda_{n}(t)=(\lambda_{n}/2)t+m_{\Lambda_{n}}(t)$, where $m_{\mathrm{M}_{n}}(t)$ and $m_{\Lambda_{n}}(t)$ are corresponding martingales (see e.g. [5, 7, 9]). Then, taking into account that $\mathsf{E}m_{\mathrm{M}_{n}}(t)=\mathsf{E}m_{\Lambda_{n}}(t)=0$, from (11) we obtain:
$0=$ $\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n+1\}\mathrm{d}\mathrm{M}_{n+1}(t)$
$+\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}2\mathsf{I}\{Z(t-)=n-2\}\mathrm{d}\Lambda_{n-2}(t)$
$-\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n\}(\mathrm{d}\mathrm{M}_{n}(t)+2\mathrm{d}\Lambda_{n}(t))$
$=$ $\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n+1\}\mu_{n+1}\mathrm{d}t$
$+\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n-2\}\lambda_{n-2}\mathrm{d}t$
$-\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n\}(\mu_{n}+\lambda_{n})\mathrm{d}t$
$=$ $\mu_{n+1}Q_{n+1}+\lambda_{n-2}Q_{n-2}-Q_{n}(\lambda_{n}+\mu_{n}).$
Relation (5) is proved.
Let us now show that the birth-and-death process $\mathscr{B}$ is recurrent if and only if the same is true for the birth-and-death process $\mathscr{A}$.
Indeed, from (3), (4) we have
$\lambda_{0}P_{0}+\sum_{n=1}^{\infty}P_{n}(\lambda_{n}+\mu_{n})=\mu_{1}P_{1}+\sum_{n=1}^{\infty}(\lambda_{n-1}P_{n-1}+\mu_{n+1}P_{n+1}),$
and from (5)–(7) we have
$\lambda_{0}Q_{0}+\sum_{n=1}^{\infty}Q_{n}(\lambda_{n}+\mu_{n})$ $=\mu_{1}Q_{1}+\mu_{2}Q_{2}+\sum_{n=2}^{\infty}(\lambda_{n-2}Q_{n-2}+\mu_{n+1}Q_{n+1})$
$=\mu_{1}Q_{1}+\sum_{n=1}^{\infty}(\lambda_{n-1}Q_{n-1}+\mu_{n+1}Q_{n+1}).$
Hence, $\sum_{n=0}^{\infty}Q_{n}=1$ if and only if $\sum_{n=0}^{\infty}P_{n}=1$.
Example 2.2.
Let $\mathscr{C}$ be a stochastic process with the same parameters $\lambda_{n}/2$ and $\mu_{n}$ as these in the birth-and-death process $\mathscr{B}$, but in any state $n$ there can be either one birth with probability $1/2$ or three births with the same probability $1/2$. It will be shown that the process $\mathscr{C}$ is recurrent if and only if the birth-and-death process $\mathscr{A}$ is recurrent.
Here we have the following system of difference equations
(12) $Q_{n}\left(\lambda_{n}+\mu_{n}\right)$ $=$ $\mu_{n+1}Q_{n+1}+\frac{\lambda_{n-1}}{4}Q_{n-1}+\frac{3\lambda_{n-3}}{4}Q_{n-3},\ n\geq 3,$
(13) $Q_{2}(\lambda_{2}+\mu_{2})$ $=$ $\mu_{3}Q_{3}+\frac{\lambda_{1}}{4}Q_{1},$
(14) $Q_{1}(\lambda_{1}+\mu_{1})$ $=$ $\mu_{2}Q_{2}+\frac{\lambda_{0}}{4}Q_{0},$
(15) $\lambda_{0}Q_{0}$ $=$ $\mu_{1}Q_{1}$
where $Q_{0}$, $Q_{1}$,…denote the final probabilities. The derivation of the system of equations (12)–(15) is similar to that given in Example 2.1 for the system of equations (5)–(7).
Similarly to the consideration given in Example 2.1, we arrive at
$\lambda_{0}Q_{0}+\sum_{n=1}^{\infty}(\lambda_{n}+\mu_{n})Q_{n}=\lambda_{0}P_{0}+\sum_{n=1}^{\infty}(\lambda_{n}+\mu_{n})P_{n},$
that means that $\sum_{n=0}^{\infty}Q_{n}=1$ if and only if $\sum_{n=0}^{\infty}P_{n}=1$. Hence, we arrive at the conclusion that the behavior of the stochastic process $\mathscr{C}$ is similar to the behavior of the birth-and-death process $\mathscr{A}$. Thus, all three processes $\mathscr{A}$, $\mathscr{B}$ and $\mathscr{C}$ are either recurrent or transient.
Apparently, the models considered in Examples 2.1 and 2.2 can be further extended, and from the point of view of recurrence or transience we will have the class of equivalent processes satisfying the same property. The class of equivalent processes should strongly contain the models of Markov chains that form connected domain only, although hitherto the provided proofs did not use any restriction. (The elementary models considered in Examples 2.1 and 2.2 do not form connected domain.)
Example 2.3.
We consider a typical example of a Markov chain that forms connected domain (call it $\mathscr{D}$), the matrix of transition probabilities of which is
$\left(\begin{matrix}0&1&0&0&0&0&0&0&0&\ldots\\ q_{1,1}&0&p_{1,1}&p_{1,2}&0&0&0&0&0&\ldots\\ q_{2,1}&q_{2,2}&0&p_{2,1}&p_{2,2}&p_{2,3}&0&0&0&\ldots\\ q_{3,1}&q_{3,2}&q_{3,3}&0&p_{3,1}&p_{3,2}&p_{3,3}&p_{3,4}&0&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right),$
where all $p_{i,j}$ and $q_{i,j}$ that appear in the matrix are assumed to be strictly positive and $\sum_{j=1}^{i}q_{i,j}+\sum_{j=1}^{i+1}p_{i,j}=1$, $i=1,2,\ldots$, and the Markov chain of the birth-and-death type (call it $\mathscr{E}$), the matrix of transition probabilities of which is
$\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ c_{1}q_{1,1}&0&c_{1}(p_{1,1}+2p_{1,2})&0&0&0&\ldots\\ 0&c_{2}(2q_{2,1}+q_{2,2})&0&c_{2}\sum_{j=1}^{3}jp_{2,j}&0&0&\ldots\\ 0&0&c_{3}\sum_{j=1}^{3}(4-j)q_{3,j}&0&c_{3}\sum_{j=1}^{4}jp_{3,j}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots\end{matrix}\right),$
where the normalization constants $c_{i}$ are given such that
$c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)=1,\quad i=1,2,\ldots.$
Our aim is to demonstrate that both Markov chains $\mathscr{D}$ and $\mathscr{E}$ are either recurrent or transient.
From the theory of birth-and-death processes, for the Markov chain $\mathscr{E}$ we have the system of equations:
(16) $c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)P_{i}$
$=c_{i-1}P_{i-1}\sum_{j=1}^{i}jp_{i-1,j}+c_{i+1}P_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j},$
(17) $P_{0}=c_{1}q_{1,1}P_{1},$
where $P_{0},P_{1},\ldots$ denote the final probabilities (if exist). In turn, for the Markov chain $\mathscr{D}$ we have the system of equation
(18) $c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)Q_{i}$
$=c_{i-1}\sum_{j=1}^{i}jp_{i-1,j}Q_{i-j}+c_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j}Q_{i+j},$
(19) $Q_{0}=c_{1}q_{1,1}Q_{1},$
where $Q_{0},Q_{1},\ldots$ denote the final probabilities (if exist). This last system of equations (18)–(19) is derived by the same method as it given in Example 2.1.
From these systems of equations (16)–(17) and (18)–(19) we respectively obtain
$P_{0}+\sum_{i=1}^{\infty}c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)P_{i}$
$=c_{1}q_{1,1}P_{1}+\sum_{i=1}^{\infty}\left(c_{i-1}P_{i-1}\sum_{j=1}^{i}jp_{i-1,j}+c_{i+1}P_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j}\right)$
and
$Q_{0}+\sum_{i=1}^{\infty}c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)Q_{i}$
$=c_{1}q_{1,1}Q_{1}+\sum_{i=1}^{\infty}\left(c_{i-1}\sum_{j=1}^{i}jp_{i-1,j}Q_{i-j}+c_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j}Q_{i+j}\right)$
$=c_{1}q_{1,1}Q_{1}+\sum_{i=1}^{\infty}\left(c_{i-1}Q_{i-1}\sum_{j=1}^{i}jp_{i-1,j}+c_{i+1}Q_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j}\right).$
Hence,
$P_{0}+\sum_{i=1}^{\infty}c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)P_{i}$
$=Q_{0}+\sum_{i=1}^{\infty}c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)Q_{i},$
and as before we arrive at the conclusion that $\sum_{n=0}^{\infty}Q_{n}=1$ if and only if $\sum_{n=0}^{\infty}P_{n}=1$.
The three examples considered above show that instead of the original Markov chain model one can study the associated birth-and-death process, if it is only required to find the condition for recurrence or transience rather than exact solution.
Let us return to our proof. The proof of the theorem does not distinguish from the proof that has already been provided for the models considered in Examples 2.12.3. Despite the model considered in Example 2.3 is particular, all the features of that proof for the general model remain unchanged. Hence, the only thing that is required to do is to consider the associated birth-and-death process and derive the required system of equations.
Denoting the birth-and-death process by $Z(t)$, we have the following system for the state probabilities at time $t+\triangle t$, where $\triangle t$ is a small increment:
(20) $\mathsf{P}\{Z(t+\triangle t)=i\}$ $=\mathsf{P}\{Z(t+\triangle t)=i~{}|~{}A_{i}^{-}(\triangle t)\}e_{i+1}^{-}\triangle t$
$+$ $\mathsf{P}\{Z(t+\triangle t)=i~{}|~{}A_{i}^{+}(\triangle t)\}e_{i-1}^{+}\triangle t$
$+$ $\mathsf{P}\{Z(t)=i\}\left[1-\triangle t\left(e_{i}^{-}+e_{i}^{+}\right)\right]+o(\triangle t),$
(21) $\mathsf{P}\{Z(t+\triangle t)=0\}$ $=\mathsf{P}\{Z(t+\triangle t)=0~{}|~{}A_{1}^{-}(\triangle t)\}e_{1}^{-}\triangle t$
$+$ $\mathsf{P}\{Z(t)=0\}(1-e_{0}^{+}\triangle t)+o(\triangle t).$
Here in (20) and (21), the meanings of the events $A_{i}^{-}(\triangle t)$ and $A_{i}^{+}(\triangle t)$ are that during time $(t,t+\triangle t]$ a down-crossing to state $i$ occurs and, respectively, an up-crossing to state $i$ occurs. The meanings of $e_{i}^{-}$ and $e_{i}^{+}$, $i\geq 1$, are the birth and, respectively, death rates in state $i$; $e_{0}^{+}=\sum_{j=1}^{\infty}jp_{0,j}$. Then the system of equations for the final probabilities of the birth-and-death process follows by the standard way. We have:
$P_{n}(e_{n}^{-}+e_{n}^{+})$ $=$ $P_{n+1}e_{n+1}^{-}+P_{n-1}e_{n-1}^{+},\quad n\geq 1,$
$P_{0}e_{0}^{+}$ $=$ $P_{1}{e_{1}^{-}}.$
Thus, the necessary and sufficient condition for recurrence of the Markov chain is $\sum_{n=1}^{\infty}\prod_{i=1}^{n}e_{i}^{-}/e_{i}^{+}=\infty.$ The theorem is proved.
Remark 2.4.
The reduction to the system of equations that describes birth-and-death processes enables us to adapt the other known results on recurrence and transience given in [1, 2] as well.
3. Example
Let us return to Example 2.3 of Section 2.3. Assuming that
$q_{i,1}=q_{i,2}=\ldots=q_{i,i}\quad\text{and}\quad p_{i,1}=p_{i,2}=\ldots=p_{i,i+1},\quad i=1,2,\ldots,$
and
$\sum_{j=1}^{i}q_{i,j}=\sum_{j=1}^{i+1}p_{i,j}=\frac{1}{2},$
we obtain the following matrix of transition probabilities:
$\left(\begin{matrix}0&1&0&0&0&0&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{1}{2}&0&\frac{1}{4}&\frac{1}{4}&0&0&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{1}{4}&\frac{1}{4}&0&\frac{1}{6}&\frac{1}{6}&\frac{1}{6}&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{1}{6}&\frac{1}{6}&\frac{1}{6}&0&\frac{1}{8}&\frac{1}{8}&\frac{1}{8}&\frac{1}{8}&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right).$
Our task is to find whether or not the Markov chain is recurrent. For associated birth-and-death process we have the following matrix of transition probabilities:
$\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{2}{5}&0&\frac{3}{5}&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&\frac{3}{7}&0&\frac{4}{7}&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&\frac{4}{9}&0&\frac{5}{9}&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right)=\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0ptq_{1}&0&p_{1}&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&q_{2}&0&p_{2}&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&q_{3}&0&p_{3}&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right),$
where the matrix on the right-hand side is added for the notational convenience.
It is readily seen that the entries of this matrix satisfy the following property: ${p_{n}}/{q_{n}}=1+(n+1)^{-1}$ for all $n\geq 1$. Recall that the birth-and-death process is recurrent if and only if $\sum_{n=1}^{\infty}\prod_{i=1}^{n}q_{i}/p_{i}=\infty$. According to Raabe’s test, if ${p_{n}}/{q_{n}}\leq 1+n^{-1}$ for all large $n$, then the aforemention series diverges. In our case it is satisfied. Hence, the birth-and-death process is recurrent. Consequently, the Markov chain is recurrent.
4. Counterexamples
The aim of this section is to show that the basic assumptions of the theorem that the states of the Markov chain form connected domain is important, and if this assumption is not satisfied the statement of the theorem fails.
Indeed, let us consider the following matrix of transition probabilities:
$\left(\begin{matrix}0&1&0&0&0&0&0&0&\ldots\\ \epsilon&0&1-\epsilon&0&0&0&0&0&\ldots\\ \frac{1}{3}-\epsilon&\epsilon&0&\epsilon&\frac{2}{3}-\epsilon&0&0&0&\ldots\\ 0&\frac{1}{2}+\epsilon&0&0&0&\frac{1}{2}-\epsilon&0&0&\ldots\\ 0&0&\frac{1}{3}-\epsilon&\epsilon&0&\epsilon&\frac{2}{3}-\epsilon&0&\ldots\\ 0&0&0&\frac{1}{2}+\epsilon&0&0&0&\frac{1}{2}-\epsilon&\ldots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\ddots\end{matrix}\right),$
where $\epsilon>0$ is a small fixed parameter. Note that the considered matrix is irreducible. However the Markov chain does not form connected domain. According to this matrix of transition probabilities, the typical excursion of the Markov chain is as follows. From state 00 it visits state $1$. Then, it visits state $2$ or otherwise goes back to the initial state. Appearing once in an even state (in our case it is state $2$), the further excursion through even states is as follows: from any state with even order number $2n$, the chain visits the state $22n+$ 2 with probability that is slightly less than $2/3$, or the state $22n-$ 2 with probability that is slightly less than $1/3$, and there is small probability to visit a state with an odd order number. An excursion over even states can be long, but once an odd state will be visited. Once visiting an odd state (distinct from $1$), the chain finally will return to state $1$, since the probability of down-crossing (which is with step 2) is higher than the such one of up-crossing. The excursion that starts from state $1$ can be repeated many times, but once the chain will return to the initial state. Our conclusion is that the Markov chain is recurrent. On the other hand, it is readily seen that in this case $\sum_{n=1}^{\infty}\prod_{i=1}^{n}{e_{i}^{-}}/{e_{i}^{+}}<\infty$. So, the statement of the theorem cannot be applied and fails.
The example above demonstrated the case of recurrent Markov chain, for which the condition of the theorem is not satisfied. It is not difficult to demonstrate a similar example for a transient Markov chain for which the condition of the theorem is not satisfied either. Here is such an example
$\left(\begin{matrix}0&1&0&0&0&0&0&0&\ldots\\ 1-\epsilon&0&\epsilon&0&0&0&0&0&\ldots\\ \frac{2}{3}-\epsilon&\epsilon&0&\epsilon&\frac{1}{3}-\epsilon&0&0&0&\ldots\\ 0&\frac{1}{2}-\epsilon&0&0&0&\frac{1}{2}+\epsilon&0&0&\ldots\\ 0&0&\frac{2}{3}-\epsilon&\epsilon&0&\epsilon&\frac{1}{3}-\epsilon&0&\ldots\\ 0&0&0&\frac{1}{2}-\epsilon&0&0&0&\frac{1}{2}+\epsilon&\ldots\\ \vdots&\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\ddots\end{matrix}\right),$
which we leave without comments because of similarity.
5. Discussion
In this paper, we established a new result for recurrence and transience for relatively wide class of Markov chains. We showed, that if the basic condition of the theorem is not satisfied, then its statement fails. The counterexamples that built in Section 3 are important. They shows that even if a Markov chain is irreducible, that is space states forms a single communicating class, there may be separate subclasses that qualitatively affect on the behavior of the Markov chain. In our case, one subclass includes all even states, another one all odd states. In these subclasses the behavior of the Markov chain demonstrated in the first example is as follows. In the even states, the state order number change rapidly up, whereas in the odd states the state order number changes slowly down. Because of specified connection between the two subclasses of states, it turns out that state $1$ is accessed with probability $1$, from which the Markov chain eventually returns to the initial state with probability $1$. In the second example, the behavior is opposite. In even states the state numbers change rapidly down, whereas in odd states the state numbers change slowly up increasing to infinity.
Data availability statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Disclosures and declarations
No conflict of interest was reported by the author
Authorship clarified
The author, who conducted this work, has no relation to any institution.